# MFSR-GAN: Multi-Frame Super-Resolution with Handheld Motion Modeling

### v1

@@ 1 Introduction
Contributions. Our primary contributions are summarized as follows. (i) We propose a realistic synthetic data engine for multi-frame super-resolution that better reflects real smartphone handheld captures by preserving sensor-specific noise and temporally correlated camera motion. (ii) We present a novel GAN-based MFSR network that emphasizes a base frame to mitigate ghosting artifacts and robustly integrates information from additional burst frames, leading to higher resolution and enhanced perceptual quality. Evaluations on both our synthetic test sets and real handheld bursts demonstrate that our method generates sharper, more detailed results compared to existing state-of-the-art MFSR methods [2, 3, 9, 10].
@@ 2 Related Work
Multi-Frame Super-Resolution. MFSR leverages information across multiple LR frames of a scene to reconstruct a HR image. Classical methods were applied to DSLR-like cameras and relied on manually derived constraints and closed-form optimizations to fuse these frames under sub-pixel misalignment for image restoration [22, 14, 11, 18]. However, estimating motion between multiple frames using these approaches is often prone to error in real-world settings, especially when dealing with non-global motion, noise, and potentially dynamic scenes.
Recent deep learning methods applied to smartphone camera images have demonstrated notable improvements in MFSR by learning data-driven representations and motion modeling. [25] demonstrated that handheld motion acquired through natural hand tremors provided sub-pixel coverage that can be exploited for super-resolution on a smartphone. [2] proposed DBSR Network by applying feature alignment and attention-based fusion on synthetic LR-HR bursts. Extending upon [2], [3] introduces additional modules for multi-frame denoising and alignment in the feature space. BIPNet [9] leveraged implicit feature alignment and pseudo-burst generation strategies for handling noisy RAW inputs. More recently, Burstormer [10] was proposed with a multi-scale design and flexible inter-frame communication to further enhance MFSR performance.
@@ 3.1 Static Short-Long Exposure Photography
To create a synthetic dataset of aligned LR-HR images, we acquired a dataset consisting of registered long exposure and short exposure MF images at HR using a mobile device following [1, 19]. All images were captured using a tripod to ensure static scenes without any global motion between frames. Utilizing a tripod eliminates unintended camera shake, allowing for perfect alignment between frames. In addition, there was no object motion in any frames. As described in Section 4, we design our model to handle object motion artifacts instead of our dataset.
@@ 3.2 Global Handheld Motion Modeling
    $\begin{bmatrix}x_{i}\\ y_{i}\\ 1\end{bmatrix}=H_{F}\begin{bmatrix}x_{1}\\ y_{1}\\ 1\end{bmatrix}\quad\text{where}\quad H_{F}=\begin{bmatrix}h_{11}&h_{12}&h_{13}\\ h_{21}&h_{22}&h_{23}\\ h_{31}&h_{32}&1\end{bmatrix}$
@@
Thus, from a dataset of real handheld captures, we sample homography matrices $H_{F}$ representing the transformation from base frame to frame F (i.e., real handheld motion), and use it to warp a sequence of short-exposure static frames, thereby inducing synthetic global motion that has the same characteristics of a handheld smartphone MF capture.
@@ 4.2 Multi-Scale Base Frame Enhancement
The multi-scale base frame enhancement module aims to align and enhance the base (reference) frame using information from the non-reference frames across multiple scales. A multi-scale approach to image restoration was first proposed by [28], and applied in a MFSR application by [10]. Similar to [10], we operate at three scales $s\in\{0,1,2\}$, where each scale corresponds to a different spatial resolution:
  • Scale 0: Original resolution, $\frac{H}{2}\times\frac{W}{2}$.
  • Scale 1: Downsampled by 2, $\frac{H}{4}\times\frac{W}{4}$.
  • Scale 2: Downsampled by 4, $\frac{H}{8}\times\frac{W}{8}$.
@@ 4.2.2 Multi-Frame Alignment & Attention
$\tilde{\mathbf{\mathfrak{F}}}_{i}^{(s)}(\mathbf{x})=\sum_{\mathbf{\delta}\in\mathcal{N}}\mathbf{w}(\mathbf{\delta})\cdot\mathbf{\mathfrak{F}}_{i}^{(s)}(\mathbf{x}+\mathbf{\delta}+\Delta\mathbf{p}_{i}^{(s)}(\mathbf{x},\mathbf{\delta})),$ (2)
@@
We also incorporate skip connections to preserve original information as depicted in Figure 3 . In addition, we introduce a ghosting skip connection, whereby we add a skip from only the base frame ($i=1$) from the RDC step to after Multi-Frame Alignment:
$\tilde{\mathbf{\mathfrak{F}}}_{i}^{(s)}=\mathbf{\mathfrak{F}}_{i}^{(s)}+\mathbf{RDC}_{1}^{(s)},\quad\text{for }i=2,\dots,F$ (4)
Combining deformable convolutions with simple attention mechanisms enables effective modeling of complex motions while focusing on relevant features, enhancing alignment quality. Skip connections with respect to base frame also allow for minimizing ghosting artifacts. Further details are in supplementary materials.
@@ 4.3 Multi-Frame Feature Enrichment
After alignment, we merge information from all frames to construct a spatially-enriched feature representation combining all frames. This module extends prior work which has found success in single frame super-resolution tasks by using long range skip connections with channel attention [29]. At Scale 0, we concatenate aligned features $\{\mathfrak{F}_{i}^{(0)}\}_{i=1}^{F}\in\mathbb{R}^{W\times H\times D}$ along the channel dimension to create a new feature representation $\mathfrak{F}_{concat}^{(0)}\in\mathbb{R}^{W\times H\times D*F}$. We treat this representation as a “single image” and apply channel attention [29] across all features of all frames. We also add a long skip connection. Different frames may provide spatio-temporal advantages and the combined spatial resolution of multiple frames will always be $\geq$ a single frame. Thus, multi-frame feature attention allows the model to access long-range context, thereby learning important spatially-enriched features across and within frames.
@@ 4.4 Image Reconstruction
The first step is to reconstruct individual frames separately. We adapt the residual-in-residual dense blocks (RRDB) from [23] for our image reconstruction. Each RRDB consists of a series of convolutional layers where each layer takes as input the concatenation of all preceding layers within the block. We stack three such layers connected via residual connections, enabling hierarchical learning of features. Given a single input frame $\mathfrak{F}_{0}^{(0)}\ \in\mathbb{R}^{W\times H\times D}$, the output of the $n$-th convolutional layer within the RRDB is computed as:
$\tilde{\mathfrak{F}}_{n}^{(0)}=\sigma\left(\mathbf{W}_{n}*\left[\mathfrak{F}_{0}^{(0)},\mathfrak{F}_{1}^{(0)},\dots,\mathfrak{F}_{n-1}^{(0)}\right]\right)+\mathfrak{F}_{0}^{(0)}$ (5)
@@
$\mathfrak{F}_{\text{RRDB}}^{(0)}=\gamma\cdot\tilde{\mathfrak{F}}_{n,3}^{(0)}\left(\tilde{\mathfrak{F}}_{n,2}^{(0)}\left(\tilde{\mathfrak{F}}_{n,1}^{(0)}\left(\mathfrak{F}_{0}^{(0)}\right)\right)\right)+\mathfrak{F}_{0}^{(0)}$ (6)
where $\tilde{\mathfrak{F}}_{n,k}^{(0)}(\cdot)$ denotes the function of the $k$-th $n$-layered residual dense block [23], and $\gamma=0.2$ is a scaling factor. This multi-level residual learning strategy allows the network to capture a wide range of features and eases the flow of gradients during training.
@@ 4.5 Discriminator
We utilize a relativistic discriminator [15] as implemented in [23]. In summary, if a standard discriminator is $D(x)=\sigma(C(x))$, a realtivistic discriminator ($D_{R}$) will be
@@ 5 Experiments
Figure 5: Qualitative comparison of our model and synthetic dataset against state-of-the-art synthetic dataset [13] and model [10].
Figure 6: Qualitative comparison of motion handling for our model against state-of-the-art [10] on real handheld MF capture. Person is waving their hands throughout MF capture. Both models trained on our synthetic dataset.
@@
As described in Section 3, we capture independent RAW short-exposure images at HR that are warped in the RGB space using a homography of real handheld multi-frame smartphone captures, thereby ensuring real spatial and temporal correlations between frames are preserved. Next, input frames are downsampled using nearest neighbor interpolation to ensure local (real) noise characteristics are also preserved. Thus, we do not add synthetic noise as our inputs already contain representative noise characteristics of a real LR input. We create two variations of our synthetic dataset: low-light and bright-light. For our low-light version, we capture 376 short-long exposure pairs, while for our bright-light dataset, we capture 107 short-long exposure pairs. All images are static, captured on a tripod using a smartphone at 50 MP (HR). The corresponding LR inputs are 12 MP. We train on 256x256 image patches. For validation, we captured thirty handheld 12 MP RAW images using the same smartphone. Validation images were captured for the purposes of photography and thus contain all distortions, artifacts and motion as would be expected in a real handheld smartphone capture.
Table 1 demonstrates that models perform relatively competitively on the Zurich synthetic dataset [13], but the overall performance is poorer for our synthetic test set. We argue this is because our synthetic dataset contains more realistic perturbations and noise, and is thus a much harder task to excel in. Figure 5 shows qualitative results of the state-of-the-art MFSR model Burstormer [10] achieving better noise handling on our dataset as compared to the Zurich dataset [13] when trained identically.

### v2

@@ 1 Introduction
Contributions. Our primary contributions are summarized as follows. (i) We propose a realistic synthetic data engine for multi-frame super-resolution that better reflects real smartphone handheld captures by preserving sensor-specific noise and temporally correlated camera motion. (ii) We present a novel GAN-based MFSR network that emphasizes a base frame to mitigate ghosting artifacts and robustly integrates information from additional burst frames using channel attention and residual connections, leading to higher resolution and enhanced perceptual quality. Evaluations on both our synthetic test sets and real handheld bursts demonstrate that our method generates sharper, more detailed results compared to existing state-of-the-art MFSR methods [2, 3, 10, 11].
@@ 2 Related Work
Multi-Frame Super-Resolution. MFSR leverages information across multiple LR frames of a scene to reconstruct a HR image. Classical methods were applied to DSLR-like cameras and relied on manually derived constraints and closed-form optimizations to fuse these frames under sub-pixel misalignment for image restoration [25, 15, 12, 20]. There has also been adjacent work done in video super-resolution that utilizes information across multiple frames [22]. However, estimating motion between multiple frames is often prone to error in real-world settings, especially when dealing with non-global motion, noise, and dynamic scenes.
Recent deep learning methods applied to smartphone camera images have demonstrated notable improvements in MFSR by learning data-driven representations and motion modeling. [29] demonstrated that handheld motion acquired through natural hand tremors provided sub-pixel coverage that can be exploited for super-resolution on a smartphone. [2] proposed DBSR Network by applying feature alignment and attention-based fusion on synthetic LR-HR bursts. Extending upon [2], [3] introduces additional modules for multi-frame denoising and alignment in the feature space. BIPNet [10] leveraged implicit feature alignment and pseudo-burst generation strategies for handling noisy RAW inputs. More recently, Burstormer [11] was proposed with a multi-scale design and flexible inter-frame communication to further enhance MFSR performance. Recently, Vision Transformers [9] have been proposed to solve various image restoration tasks, but they are difficult to implement on mobile devices [27] and ConvNets have been shown to outperform them on smaller datasets [9, 17].
@@ 3.1 Static Short-Long Exposure Photography
To create a synthetic dataset of aligned LR-HR images, we acquired a dataset consisting of registered long exposure and short exposure MF images at HR using a mobile device following [1, 21]. All images were captured using a tripod to ensure static scenes without any global motion between frames. Utilizing a tripod eliminates unintended camera shake, allowing for perfect alignment between frames.
@@ 3.2 Global Handheld Motion Modeling
    $\begin{bmatrix}x_{i}\\ y_{i}\\ 1\end{bmatrix}=H_{F}\begin{bmatrix}x_{1}\\ y_{1}\\ 1\end{bmatrix},\quad H_{F}=\begin{bmatrix}h_{11}&h_{12}&h_{13}\\ h_{21}&h_{22}&h_{23}\\ h_{31}&h_{32}&1\end{bmatrix}$ (1)
@@
Thus, from a dataset of real handheld captures, we sample homography matrices $H_{F}$ representing the transformation from base frame to frame F (i.e., real handheld motion), and use it to warp a sequence of short-exposure static frames, thereby inducing synthetic global motion that has the same characteristics of a handheld smartphone MF capture. Since the matrices are sampled independently, temporal correlations are not captured in this model because it would require extensive data collection in order to generate independent groups of homographies. We instead rely on modelling the marginal probability distributions of homographies of true handheld motion, but we note the importance of joint distributions for future work.
@@ 4.2 Multi-Scale Base Frame Enhancement
The multi-scale base frame enhancement module aims to align and enhance the base (reference) frame using information from the non-reference frames across multiple scales. A multi-scale approach to image restoration was first proposed by [32], and applied in a MFSR application by [11]. Similar to [11], we operate at three scales $s\in\{0,1,2\}$, where each scale corresponds to a different spatial resolution.
@@ 4.2.2 Multi-Frame Alignment & Attention
$\tilde{\mathbf{\mathfrak{F}}}_{i}^{(s)}(\mathbf{x})=\sum_{\mathbf{\delta}\in\mathcal{N}}\mathbf{w}(\mathbf{\delta})\cdot\mathbf{RDC}_{i}^{(s)}(\mathbf{x}+\mathbf{\delta}+\Delta\mathbf{p}_{i}^{(s)}(\mathbf{x},\mathbf{\delta})),$ (3)
@@
We also incorporate skip connections to preserve original information as depicted in Figure 3 . In addition, we introduce a ghosting skip connection, whereby we add a skip from only the base frame feature ($i=1$) to after Multi-Frame Alignment:
$\hat{\mathbf{\mathfrak{{F}}}}_{i}^{(s)}=\mathbf{\mathfrak{\tilde{F}}}_{i}^{(s)}+\mathbf{\mathfrak{F}}_{1}^{(s)},\quad\text{for }i=2,\dots,F$ (5)
Combining deformable convolutions with simple attention mechanisms enables effective modeling of complex motions while focusing on relevant features, enhancing alignment quality. Skip connections with respect to base frame also allow for minimizing ghosting artifacts.
@@ 4.3 Multi-Frame Feature Enrichment
After alignment, we merge information from all frames to construct a spatially-enriched feature representation combining all frames. This module extends prior work which has found success in single frame super-resolution tasks by using long range skip connections with channel attention [33]. At Scale 0, we concatenate aligned features $\{\mathfrak{\hat{F}}_{i}^{(0)}\}_{i=1}^{F}\in\mathbb{R}^{W\times H\times D}$ along the channel dimension to create a new feature representation $\mathfrak{\hat{F}}_{concat}^{(0)}\in\mathbb{R}^{W\times H\times D*F}$. We treat this representation as a “single image” and apply channel attention [33] across all features of all frames. We also add a long skip connection. Different frames may provide spatio-temporal advantages and the combined spatial resolution of multiple frames will always be $\geq$ a single frame. Thus, multi-frame feature attention allows the model to access long-range context, thereby learning important spatially-enriched features across and within frames.
@@ 4.4 Image Reconstruction
The first step is to reconstruct individual frames separately. We adapt the residual-in-residual dense blocks (RRDB) from [26] for our image reconstruction. Each RRDB consists of a series of convolutional layers where each layer takes as input the concatenation of all preceding layers within the block. We stack three such layers connected via residual connections, enabling hierarchical learning of features. Given a single input frame $\mathfrak{\hat{F}}_{0}^{(0)}\ \in\mathbb{R}^{W\times H\times D}$, the output of the $n$-th convolutional layer within the RRDB is computed as:
~$\tilde{\mathfrak{R}}_{n}^{(0)}=\sigma\left(\mathbf{W}_{n}*\left[\mathfrak{\hat{F}}_{0}^{(0)},\mathfrak{\hat{F}}_{1}^{(0)},\dots,\mathfrak{\hat{F}}_{n-1}^{(0)}\right]\right)+\mathfrak{\hat{F}}_{0}^{(0)}$ (6)
@@
$\mathfrak{R}_{\text{RRDB}}^{(0)}=\gamma\cdot\tilde{\mathfrak{R}}_{n,3}^{(0)}\left(\tilde{\mathfrak{R}}_{n,2}^{(0)}\left(\tilde{\mathfrak{R}}_{n,1}^{(0)}\left(\mathfrak{R}_{0}^{(0)}\right)\right)\right)+\mathfrak{R}_{0}^{(0)}$ (7)
where ~$\tilde{\mathfrak{R}}_{n,k}^{(0)}(\cdot)$ denotes the function of the $k$-th $n$-layered residual dense block [26], and $\gamma=0.2$ is a scaling factor. This multi-level residual learning strategy allows the network to capture a wide range of features and eases the flow of gradients during training.
@@ 4.5 Discriminator
We utilize a relativistic discriminator [16] as implemented in [26]. In summary, if a standard discriminator is $D(x)=\sigma(C(x))$, a relativistic discriminator ($D_{R}$) will be
@@ 5 Experiments
Figure 5: Qualitative comparison of our model and synthetic dataset against SOTA synthetic dataset [14] and model [11].
Figure 6: Qualitative comparison of motion handling for our model against Burstormer [11] on real handheld MF capture. Person is waving their hands throughout MF capture. Both models trained on our synthetic dataset.
@@
As described in Section 3, we capture independent RAW short-exposure images at HR that are warped in the RGB space using a homography of real handheld multi-frame smartphone captures, thereby ensuring real spatial and temporal correlations between frames are preserved. Next, input frames are downsampled using nearest neighbor interpolation to ensure local (real) noise characteristics are also preserved. Thus, we do not add synthetic noise as our inputs already contain representative noise characteristics of a real LR input. We create two variations of our synthetic dataset: low-light and bright-light. For our low-light version, we capture 376 short-long exposure pairs, while for our bright-light dataset, we capture 107 short-long exposure pairs. All images are static, captured on a tripod using a smartphone at 50 MP (HR). The corresponding LR inputs are 12 MP. We train on 256x256 image patches. For validation, we captured thirty handheld 12 MP RAW images using the same smartphone. Validation images were captured for the purposes of photography and thus contain all distortions, artifacts and motion as would be expected in a real handheld smartphone capture. Since these images represent “real” handheld photography, one-to-one real ground truth images cannot be captured and qualitative evaluation is performed as shown in Figures 4 and 6.
Table 1 demonstrates that models perform relatively competitively on the Zurich synthetic dataset [14], but the overall performance is poorer for our synthetic test set. We argue this is because our synthetic dataset contains more realistic perturbations and noise, and is thus a much harder task to excel in. We have added more results in supplementary material showing the importance of accurate motion modeling. Figure 5 shows qualitative results of the state-of-the-art MFSR model Burstormer [11] achieving better noise handling on our dataset as compared to the Zurich dataset [14] when trained identically.
@@ 6 Conclusion
7 Synthetic Data Engine
We provides more details of our synthetic data egnine. Figure 7 provides more examples of the temporal and spatial correlation of motion within a multi-frame data capture. Figure 8 shows visual results for the SOTA model in MFSR [10] trained on our dataset with motion sampled from a uniform distribution of rotations and translations against sampling from a dataset of real handheld homography matrices.
@@ 7 Synthetic Data Engine
Figure 8: Burstormer [11] trained on our dataset with motion vectors sampled from a uniform distribution of pre-defined rotations and translation parameters (left) versus Burstormer trained on the same dataset but with motion vectors sampled from a dataset of homography matrices defining the transformations within a real handheld burst photography capture as per our synthetic data engine (right). All experimental conditions identical.
8 Model Architecture Details
We provide more details for our model architecture previously introduced in Section 4. Our model architecture builds upon the work previously completed for MFSR by [11]. Our adaptions of the Multi-Frame Alignment and Multi-Frame Attention modules are presented in Figure 9 with descriptions in Section 4.2.2.